OpenStudy (anonymous):

Find the complex fifth root of 1. I don't even know where to begin? haha.. I am guessing it looks like this sqrt(1) (0 + I)^5 Can someone help me on this please?

OpenStudy (zugzwang):

You know how to take roots of complex numbers? Say they're of the form \[z = r(\cos \theta + i \sin \theta)\]

OpenStudy (anonymous):

aw, yes. DeMoivre's Theorem. but I am still confused on how to solve it

OpenStudy (anonymous):